OpenStudy (anonymous):

13. A transformer has a primary coil with 600 turns and a secondary coil with 300 turns. If the output voltage is 320 volts, what is the input voltage? (1 point)

OpenStudy (shane_b):

For transformers the voltage ratio is equal to the turn ratio: \[\frac{Vs}{Vp}=\frac{Ns}{Np}\]so plugging in values you get:\[\frac{320V}{V_p}=\frac{300}{600}\]\[V_p=640V\]